Science Operations

Instrument Status Post SSD 2022

Posted Sep 19, 2022



The following summarizes the work done to the facility instruments during summer shutdown 2022, and the status of those instruments post restart:


Both LBCs have been checked out and are operating nominally. During the summer shutdown, the filters and the accessible lenses (L1 front and back & the front of L2) were cleaned.  As far as the LBC operation, no changes have been noted since last semester. There is a flaw in the guiding code which may affect guiding non-sidereal targets where the rate x exposure time is greater than ~2″ and an image jump may be seen. We are working to release an updated version of the LBC SW that will prevent spurious guide corrections; after that, NS targets will be guided until their rate x exposure time is ~2″, after which they will not be guided, but no image jumps should be seen.
The TMS (telescope metrology system) fibers, collimators and retroreflectors were all installed and the system was checked out closed-dome, however weather and time did not cooperate to complete a full on-sky checkout of the system during the engineering nights. It has since been used for one half-night, and no issues were seen. The intention is to continue to use the TMS system with the LBCs during partner science nights in a shared risk capacity, as was done last semester.


The G150 grating was removed from LUCI1 this summer and replaced by the G040 grating, making the grating configurations in the 2 LUCI’s identical.  The release of the G040 grating is pending commissioning.  While LUCI1 was off telescope, the entrance window was cleaned, removing the internal debris noted last semester. Maintenance and internal inspections were completed, including a detailed inspection of the detector focus mechanism.  LUCI2 remained on telescope.

All standard tests of LUCI1 and LUCI2 were completed after the summer shutdown. The new transforms and field aberration calculations were all confirmed.

All science modes for LUCI1 and LUCI2 have been calibrated and verified closed dome including:

Seeing limited imaging and G210/N1.8 spectroscopy were verified on sky, and typical science cases with G200/N1.8 spectroscopy and N30 imaging are pending on sky verification.  All modes are released with a note that additional sky time is desired for further verification.


The SX Adsec was removed for the repair of 10 detached magnets, restoring full functionality at the edge.  DX Adsec was also briefly removed for some critical maintenance.

The DX Adsec was checked out fully for seeing limited operations.  LUCI AO or TTM is not an available mode at this time, and still in the commissioning phase.

The SX Adsec was recalibrated for seeing limited operations.  The diffraction limited and ESM operations were verified closed-dome for all binnings.  On-sky checkouts of bin1 and bin2 were completed during the restart but this was prior to the generation of updated NCPA tables.  New tables have been generated and are in the process of verification.  All diffraction limited and ESM modes are released for science in shared risk until additional on-sky tests are completed.


MODS1 and MODS2 were installed on Sept 20, 2022. All of the mechanisms have been exercised and are working well. Images have been taken with the guide and WFS cameras, and over a hundred biases and calibration test images have been taken closed-dome.

On Sep 25 UT, we were able to perform the on-sky checkout, checking the AGw alignment and verifying that the rotator zero point, which aligns the slits and masks along the N/S axis for PA=0, has not changed, so the AGw stage-to-sky transforms from last year are still usable. The night was clear and we obtained data on a spectrophotometric standard. The instruments are performing well, though we have noted a few bad readouts with MODS2B (4 of ~250 images) and continue to see intermittent readout delays for both MODS1R and, now also, MODS2B.

During the summer, a team from OSU (Mark Derwent, Dan Pappalardo, and Mike Engelman) came to the LBT to work on repairing the slitmask cassettes, both of which had sustained damage in the past year, and on preparing ready-to-go spares that will make replacement of the dichroic, filter wheel,and grating wheel pinion gears, if needed, faster and easier. They are adding steel sensor detection targets to the mask frames which will enable a larger sensor gap and make the mechanism more robust against the type of failure experienced last year. The slitmask cassette in MODS2 can once again be loaded to capacity, though in MODS1 a bent loading chute is effectively removing 2 custom positions until it can be fixed. The optics were inspected and dusted off before the instrument was mounted. Flat field images taken show no significant debris except for a spot on the MODS1 dichroic, but this has a special multi-layer coating and cleaning is on hold until we settle on a safe way to do it.


The SX primary was realuminized and DX primary washed this year.  The SX m3 was recoated this year as well.